Midnight97:

A 50-kilogram student is running and has 225 joules of kinetic energy. The student’s speed is____ meters/second.

flowerpower52:

The student’s speed is meters/second ✓ 225 = 1/2 (50) (v2)225 = 25 (v2)225/25 = v29 = v2√9 = vv = 3 m/s